Hi I'm having trouble answering this question
'In a triangle,the largest angle is 20 degrees larger than the smallest angle.Write down the three angles it could be?'
The triangle I've got is that one of the side of the triangle is the largest,the other side is in the middle which is the second largest and the last one is the smallest side.
I've tried different ways like starting from 180 degrees and took away from there but they don't actually work and I don't get it.This is a question from my practice test for revision and it's a calculator test.
Thanks from Naomi

Answer by Alan3354(69443)

In a triangle,the largest angle is 20 degrees larger than the smallest angle.Write down the three angles it could be?'
---------
Call them A, B & C
A = B + 20
A + B + C = 180
C = 180 - (A + B)
C = 180 - (2B + 20)
C = 160 - 2B
---------
B < C < A since A is the largest and B is the smallest
B < 160 - 2B
3B < 160
B < 160/3 degs
--> A < 220/3 degs
A + B < 380/3
----
More later today